OpenStudy (anonymous):

t = (14-3r)/(6) for r

mathslover (mathslover):

first of all : \[\Huge{\mathbb{WElcome}\textbf{To}\mathbb{OPen}\textbf{Study}}\]

mathslover (mathslover):

\[\huge{t=\frac{14-3r}{6}}\] \[\huge{6t=14-3r}\]

OpenStudy (anonymous):

Thans (:

OpenStudy (anonymous):

Thanks*